# One Stone with Two Birds: A Null-Text-Null Frequency-Aware Diffusion Models for Text-Guided Image Inpainting

### v1

@@ 1 Introduction
So far, the prior arts still fail to simultaneously overcome the above two challenges. Such failure, as investigated in Fig.2, stems from the following hurdle: as disclosed by qian2024boosting , the low-frequency band fluctuates more significantly during the early stage of the denoising process with high-level noise than the mid-and-high frequency bands, since diffusion models typically recover the low-frequency band first and progressively refine the mid-and-high frequency bands [^1] Compared to the mid-frequency band, the high-frequency band is sparser and more easy to be disrupted by high-level noise in the early stages, resulting in much less contribution on guiding the low-frequency band than mid-frequency band. Therefore, we focus on mid-and-low frequency bands.. Therefore, the low-frequency band for both masked and unmasked regions are easy to be modulated by text prompts, as illustrated in Fig.2(a). To be contrary, as seen in Fig.2(b), the mid-frequency band across all regions is robust to the text prompts while aligns well with text prompts, which may better preserve the unmasked regions than low-frequency band upon text prompts. The above observations further motivate us to decompose the semantics consistency across masked and unmasked regions into the individual task of achieving that as per each frequency band across masked and unmasked regions, while preserving the unmasked regions, to tackle two challenges in a row. To this end, we formally delve into the following questions: How to disentangle different frequency bands, particularly the early stage of the denoising process with high-level noise? and How to exploit the hybrid frequency bands for diffusion models to simultaneously achieve the above two goals?
@@ 2.2 Motivation: Hybrid Frequency Aware Diffusion Models to Text-Guided Image Inpainting
Capitalizing on the discoveries, instead of low-frequency, we propose to exploit the mid-frequency band during denoising process, which plays the pivotal role of achieving the semantics consistency upon text prompts, while can better achieve the semantics consistency between masked and unmasked regions, while preserve its own frequency band well, than low-frequency band, owing to its robustness to the text prompt during the denoising process. The above observations further motivate us with a novel frequency-aware null-text-null diffusion models, named NTN-Diff, for text-guided image inpainting. Our pipeline is illustrated in Fig.3, where we discuss NTN-Diff within the early and late stages, demarcated by the critical step $\lambda T$ with T as the total steps for diffusion models; the detailed discussions for $\lambda$ can be seen in Sec.3.3.2. (See more intuitions on hybrid frequency aware diffusion models to text-guided image inpainting in Sec.A of the Appendix)
@@ 2.4 Late Stage of Text-Guided Denoising Process
where $z^{gt}_{T-t}$ donates the forward diffused results with the (Tt)(T-t)-th step. Building upon the noisy sample $\hat{z}^{mid}_{t}$, the text prompt $c$, along with the step $t$, we perform the denoising network $\epsilon_{\theta}(\hat{z}^{mid}_{t},t,\tau_{\theta}(c))$ to get the denoising result ${z}^{mid}_{t-1}$ in the latent space for the next (t1)(t-1)-th step. After $\lambda T$ steps, the denoised latent representation ${z}^{mid}_{0}$ is reconstructed into the pixel space via the VAE decoder $\mathcal{D}$ to generate the final image inpainting output.
@@ 3.2 Comparison with State-of-the-arts
Quantitative and Qualitative analysis. The quantitative results in Table.1 summarizes our findings below: NTN-Diff enjoys larger IR, HPS v2, PSNR and CLIP Score, together with smaller MSE and LPIPS than the competitors. Notably, BrushNet remains the large performance margins (at most 0.73% for IR, 0.62% for HPS, 8.59% for PSNR, 30.18% for MSE, 15.51% for LPIPS and 0.23% for CLIP Score) compared to NTN-Diff in Table.1. We also present the quantitative results of NTN-Diff with the pixel-level blending operation of ju2024brushnet , named NTN-Diff*, to preserve the unmasked regions, which demonstrates the ability to tame the hybrid frequency issue, the results further verifies the intuition in Sec.1 – NTN-Diff achieves the semantics consistency between mid-and-low frequency bands across masked and unmasked regions, while preserving unmasked regions. We illustrate the above intuitions in Fig.5. (More compared results for Editbench wang2023imagen can be seen in Sec.B of the Appendix).
@@ 3.3.1 Discussion on the Different Denoising Process of the Early Stage
To validate of our NTN-Diff(Sec.2.3), we perform the ablation study on BrushBench and EditBench datasets with three variants: Case A: removing masked self-attention in the first null-text denoising process (Sec.2.3.1); Case B: removing both null-text and text-guided denoising process (Sec.2.3.2). Case C: removing the null-text denoising process (Sec.2.3.3); Table.3 reports the results for all three cases, which suggests that NTN-Diff outperforms Case A, confirming that the null-text denoising process conditioned on null-text prompt can avoid being influenced by text prompts even under the high-level noise, focusing primarily on low-frequency band, which in line with Sec.2.3.1; Case B exhibits a suboptimal performance degradation, despite the null-text denoising process in the early stage, together with the text-guided denoising process, which verifies that the null-text mid-frequency guided denoised process can denoise the low-frequency band for the masked regions to be aligned with text prompt, owing to the denoised mid-frequency from text-guided process, hence validating the important of text-guided process (Sec.2.3.2); NTN-Diff maintains achieves the large performance gain over Case C (at most 44.8%), implying that only mid-frequency aligned with text prompts for masked and unmasked regions, still failed to achieve the consistency across the whole regions, which is in line with Sec.2.3.3 and illustrated in Fig.6. (See more generation results for the impact of the text and null-text prompts on low-and-mid frequency bands in Sec.C of the Appendix).
@@ 3.3.2 Hyperparameter Sensitivity Analysis of $\lambda$ for the Length of the Early and Late Stage
We analyze the parameter $\lambda$ (Sec.2.2) that divides the denoising process into early and late stages at step $\lambda T$, we set the $\lambda$ values to {0.9; 0.8; 0.7; 0.6; 0.5}. As reported in Table.3, when $\lambda=0.9$, the performance is the worst with the shortest early stage. When $\lambda=0.6$ with the balance of early and late stage, the best performance are achieved, confirming that the desirable length of early stage can make the low-frequency band under the mid-frequency guidance for masked regions achieve the consistency to the substituted unmasked regions for ground truth, which is consistent to Sec.2.4. (See more generation results in Sec.D of the Appendix).
@@ 80. Guidelines:
  • More intuitions on hybrid frequency aware diffusion models to text-guided image inpainting, as mentioned in Sec.2.2 of the main body (Sec.A)
  • Additional quantitative results and qualitative analysis with higher resolution for the comparison with state-of-the-arts, as mentioned in Sec.3.2 of the main body (Sec.B).
  • More ablation studies on the impact of text and null-text prompts on low-and-mid frequency bands during denoising process, as mentioned in Sec.3.3.1 of the main body (Sec.C).
  • Additional generation results for the ablation study about hyperparameter sensitivity analysis of $\lambda$ for the length of the early and late stage, as mentioned in Sec.3.3.2 of the main body (Sec.D).
  • We list the limitations and broader impacts (Sec.E).
Appendix A More Intuition on Hybrid Frequency Aware Diffusion Models to Text-Guided Image Inpainting
@@
Appendix B Additional Quantitative and Qualitative analysis in the Sec.3.2 of the Main body
@@
Appendix C More Ablation Studies on the Impact of Text and Null-Text Prompts on Low-and-Mid Frequency Bands during Denoising Process
@@
Appendix D Additional Generation Results for the Ablation Study in Sec.3.3.2 of the Main Body
@@
Appendix E Limitations and Broader Impacts

### v2

@@ 1 Introduction
So far, the prior arts still fail to simultaneously overcome the above two challenges. Such failure, as investigated in Fig.2, stems from the following hurdle: as disclosed by qian2024boosting , the low-frequency band fluctuates more significantly during the early stage of the denoising process with high-level noise than the mid-and-high frequency bands, since diffusion models typically recover the low-frequency band first and progressively refine the mid-and-high frequency bands [^1] Compared to the mid-frequency band, the high-frequency band is sparser and more easy to be disrupted by high-level noise in the early stages, resulting in much less contribution on guiding the low-frequency band than mid-frequency band. Therefore, we focus on mid-and-low frequency bands.(See more results in Sec.A of the Appendix). Therefore, the low-frequency band for both masked and unmasked regions are easy to be modulated by text prompts, as illustrated in Fig.2(a). To be contrary, as seen in Fig.2(b), the mid-frequency band across all regions is robust to the text prompts while aligns well with text prompts, which may better preserve the unmasked regions than low-frequency band upon text prompts. The above observations further motivate us to decompose the semantics consistency across masked and unmasked regions into the individual task of achieving that as per each frequency band across masked and unmasked regions, while preserving the unmasked regions, to tackle two challenges in a row. To this end, we formally delve into the following questions: How to disentangle different frequency bands, particularly the early stage of the denoising process with high-level noise? and How to exploit the hybrid frequency bands for diffusion models to simultaneously achieve the above two goals?
@@ 2.2 Motivation: Hybrid Frequency Aware Diffusion Models to Text-Guided Image Inpainting
Capitalizing on the discoveries, instead of low-frequency, we propose to exploit the mid-frequency band during denoising process, which plays the pivotal role of achieving the semantics consistency upon text prompts, while can better achieve the semantics consistency between masked and unmasked regions, while preserve its own frequency band well, than low-frequency band, owing to its robustness to the text prompt during the denoising process. The above observations further motivate us with a novel frequency-aware null-text-null diffusion models, named NTN-Diff, for text-guided image inpainting. Our pipeline is illustrated in Fig.3, where we discuss NTN-Diff within the early and late stages, demarcated by the critical step $\lambda T$ with T as the total steps for diffusion models; the detailed discussions for $\lambda$ can be seen in Sec.3.3.2. (See more intuitions on hybrid frequency aware diffusion models to text-guided image inpainting in Sec.B of the Appendix)
@@ 2.4 Late Stage of Text-Guided Denoising Process
where $z^{gt}_{T-t}$ donates the forward diffused results with the (Tt)(T-t)-th step. Building upon the noisy sample $\hat{z}^{mid}_{t}$, the text prompt $c$, along with the step $t$, we perform the denoising network $\epsilon_{\theta}(\hat{z}^{mid}_{t},t,\tau_{\theta}(c))$ to get the denoising result ${z}^{mid}_{t-1}$ in the latent space for the next (t1)(t-1)-th step. After $\lambda T$ steps, the denoised latent representation ${z}^{mid}_{0}$ is reconstructed into the pixel space via the VAE decoder $\mathcal{D}$ to generate the final image inpainting output. (See the pseudo algorithm in Sec.C of the Appendix)
@@ 3.2 Comparison with State-of-the-arts
Quantitative and Qualitative analysis. The quantitative results in Table.1 summarizes our findings below: NTN-Diff enjoys larger IR, HPS v2, PSNR and CLIP Score, together with smaller MSE and LPIPS than the competitors. Notably, BrushNet remains the large performance margins (at most 0.73% for IR, 0.62% for HPS, 8.59% for PSNR, 30.18% for MSE, 15.51% for LPIPS and 0.23% for CLIP Score) compared to NTN-Diff in Table.1. We also present the quantitative results of NTN-Diff with the pixel-level blending operation of ju2024brushnet , named NTN-Diff*, to preserve the unmasked regions, which demonstrates the ability to tame the hybrid frequency issue, the results further verifies the intuition in Sec.1 – NTN-Diff achieves the semantics consistency between mid-and-low frequency bands across masked and unmasked regions, while preserving unmasked regions. We illustrate the above intuitions in Fig.5. (See more analysis of computational efficiency in Sec.D of the Appendix, along with more compared results for Editbench wang2023imagen can be seen in Sec.E of the Appendix)
@@ 3.3.1 Discussion on the Different Denoising Process of the Early Stage
To validate of our NTN-Diff(Sec.2.3), we perform the ablation study on BrushBench and EditBench datasets with three variants: Case A: removing masked self-attention in the first null-text denoising process (Sec.2.3.1); Case B: removing both null-text and text-guided denoising process (Sec.2.3.2). Case C: removing the null-text denoising process (Sec.2.3.3); Table.3 reports the results for all three cases, which suggests that NTN-Diff outperforms Case A, confirming that the null-text denoising process conditioned on null-text prompt can avoid being influenced by text prompts even under the high-level noise, focusing primarily on low-frequency band, which in line with Sec.2.3.1; Case B exhibits a suboptimal performance degradation, despite the null-text denoising process in the early stage, together with the text-guided denoising process, which verifies that the null-text mid-frequency guided denoised process can denoise the low-frequency band for the masked regions to be aligned with text prompt, owing to the denoised mid-frequency from text-guided process, hence validating the important of text-guided process (Sec.2.3.2); NTN-Diff maintains achieves the large performance gain over Case C (at most 44.8%), implying that only mid-frequency aligned with text prompts for masked and unmasked regions, still failed to achieve the consistency across the whole regions, which is in line with Sec.2.3.3 and illustrated in Fig.6. (See more generation results for the impact of the text and null-text prompts on low-and-mid frequency bands in Sec.F of the Appendix).
@@ 3.3.2 Hyperparameter Sensitivity Analysis of $\lambda$ for the Length of the Early and Late Stage
We analyze the parameter $\lambda$ (Sec.2.2) that divides the denoising process into early and late stages at step $\lambda T$, we set the $\lambda$ values to {0.9; 0.8; 0.7; 0.6; 0.5}. As reported in Table.3, when $\lambda=0.9$, the performance is the worst with the shortest early stage. When $\lambda=0.6$ with the balance of early and late stage, the best performance are achieved, confirming that the desirable length of early stage can make the low-frequency band under the mid-frequency guidance for masked regions achieve the consistency to the substituted unmasked regions for ground truth, which is consistent to Sec.2.4. (See more generation results in Sec.G of the Appendix).
@@ 80. Guidelines:
  • More analysis of the high-and-low frequency components in the denoising process, as mentioned in Sec.1 of the mainbody (Sec.A)
  • More intuitions on hybrid frequency aware diffusion models to text-guided image inpainting, as mentioned in Sec.2.2 of the mainbody (Sec.B)
  • The algorithm of NTN-Diff, as mentioned in Sec.2.4 of the mainbody (Sec.C)
  • More analysis of computational efficiency for the comparison with state-of-the-arts, as mentioned in Sec.3.2 of the mainbody (Sec.D).
  • Additional quantitative results and qualitative analysis with higher resolution for the comparison with state-of-the-arts, as mentioned in Sec.3.2 of the mainbody (Sec.E).
  • More ablation studies on the impact of text and null-text prompts on low-and-mid frequency bands during denoising process, as mentioned in Sec.3.3.1 of the mainbody (Sec.F).
  • Additional generation results for the ablation study about hyperparameter sensitivity analysis of $\lambda$ for the length of the early and late stage, as mentioned in Sec.3.3.2 of the mainbody (Sec.G).
  • We list the limitations and broader impacts (Sec.H).
Appendix A More Analysis of the High-and-Low Frequency Components in the Denoising Process
Due to the page limitation, we further conduct experiments to verify the conclusion stated in Sec. 1 of the mainbody: “the low-frequency band fluctuates more significantly during the early stage of the denoising process with high-level noise than the mid- and high-frequency bands.” This observation aligns with the common understanding that diffusion models first recover low-frequency components and progressively refine mid- and high-frequency details. Motivated by this, we experimentally validate the denoising process using several metrics to measure variations in the low-and-high frequency bands at different timesteps, as reported in Table. 4.
We adopt two groups of metrics to analyze the frequency behavior during the denoising process:
  • Low-frequency information: We adopt Mean Gray, Mean HSV-V, and Mean Luma, which primarily capture the global brightness and luminance structure of the denoised image. These metrics evaluate how well the low-frequency bands are maintained throughout the denoising process.
  • High-frequency information: We use Average Gradient, Variance, and LBP Variance, which are sensitive to edge sharpness, local contrast, and texture complexity, respectively. These metrics reflect the preservation and recovery of fine details during denoising.
The results show that low-frequency metrics (Mean Gray, HSV-V, Luma) remain nearly unchanged from the beginning, indicating early stabilization. In contrast, high-frequency metrics (Avg Gradient, Variance, LBP Variance) increase steadily in the later timesteps, confirming that high-frequency details are refined gradually during denoising, which is consistent with the statement in Sec. 1 of the mainbody.
Table 4: Metrics measuring the low-and-high frequency band variations during the denoising process, confirming that diffusion models typically recover the low-frequency band first and progressively refine the mid-and-high frequency bands.
Timesteps Mean Gray Mean HSV-V Mean Luma Avg Gradient Variance LBP Variance
0 120.9 136.0 120.9 59.01 3955.0 4.45
5 120.8 134.6 120.8 75.07 4098.3 3.71
10 120.7 133.5 120.7 101.58 4378.9 3.25
15 120.7 133.7 120.7 133.26 4790.8 3.07
20 120.9 135.1 120.9 164.23 5274.8 3.07
25 121.2 137.3 121.2 190.49 5743.9 3.15
30 121.4 139.2 121.4 209.84 6099.5 3.25
35 121.1 140.0 121.1 222.03 6291.1 3.32
40 120.5 140.1 120.5 228.75 6363.9 3.35
45 119.9 139.8 119.9 232.19 6394.8 3.38
50 119.5 139.5 119.5 233.88 6419.5 3.40
Appendix B More Intuition on Hybrid Frequency Aware Diffusion Models to Text-Guided Image Inpainting
@@
Appendix C The Algorithm of NTN-Diff
Our proposed NTN-Diff pipeline consists of a null-text denoising process (Sec. 2.3.1) that avoids the influence of text prompts, and a text-guided denoising process (Sec. 2.3.2) that reconstructs the masked regions while replacing the low-frequency band of its denoised output with that from the null-text denoising process. Building upon this design, we further utilize the denoised mid-frequency information to guide another null-text denoising process (Sec. 2.3.3) by substituting the mid-frequency band accordingly. In addition, a late-stage text-guided denoising process (Sec. 2.4) is conducted, during which the unmasked regions from the early stage of the diffusion process are progressively substituted to preserve structural consistency. Based on the above, we summarize the complete algorithm of our proposed null-text-null frequency-aware diffusion model and present the detailed steps in Algorithm 1 as follow:
Algorithm 1 NTN-Diff
Input: masked image $I$, binary mask $M$, text prompt $C$, null-text prompt $C_{\emptyset}$
Output: inpainted image $I^{\prime}$
I. Diffusion Process for Unmasked Regions
1:Extract the initial latent feature of the unmasked region $z_{0}^{gt}=E(I)$.
2:for $t=0$ to $T_{inv}-1$ do
3:  Compute $z_{t+1}^{gt}$ from $z_{t}^{gt}$;
4:end for{DDIM inversion}
II. Early Stage
1:Initialize $z_{T}^{un}$, $z_{T}^{text}$, $z_{T}^{mid}\sim\mathcal{N}(0,1)$.
2:for $t=T$ to $\lambda T+1$ do
3:  Substitute the unmasked regions of $z_{t}^{un}$ with the corresponding counterpart of $z_{t}^{gt}$ via Eq.3;
4:  Compute $z_{t-1}^{un}$ from $z_{t}^{un}$, conditioned on $C_{\emptyset}$ via Eq.4;
5:  Substitute low-frequency band of $z_{t}^{text}$ with the corresponding counterpart of $z_{t}^{un}$ via Eq.5–7;
6:  Compute $z_{t-1}^{text}$ from $z_{t}^{text}$, conditioned on $C$;
7:  Substitute mid-frequency band of $z_{t}^{mid}$ with the corresponding counterpart of $z_{t}^{text}$ via Eq.8–10;
8:  Compute $z_{t-1}^{mid}$ from $z_{t}^{mid}$, conditioned on $C_{\emptyset}$;
9:end for
III. Late Stage
1:for $t=\lambda T$ to 11 do
2:  Substitute unmasked regions of $z_{t}^{mid}$ with the corresponding counterpart of $z_{t}^{gt}$ via Eq.11;
3:  Compute $z_{t-1}^{mid}$ from $z_{t}^{mid}$, conditioned on $C$;
4:end for
5:Output the final inpainted image $I^{\prime}=D(z_{0}^{mid})$.
Appendix D More Analysis of Computational Efficiency in the Sec.3.2 of the Mainbody
As indiacted in the mainbody, we further provide the inference time and VRAM consumption of our proposed NTN-Diff model, and compared it with several representative baselines, which are reported in Table. 5. Our inference time is comparable to BrushNet ju2024brushnet and faster than recent methods such as HDP manukyan2023hd and PP zhuang2023task . Moreover, our model consumes less VRAM than other text-guided inpainting models, owing to the fact that our method performs low- and mid-frequency substitution across different branches after each denoising step. The key idea is that, being decoupled from the resource-intensive denoising process itself, our approach introduces only negligible additional memory cost compared to the single denoising process in BLD avrahami2023blended .
Table 5: Inference time and VRAM consumption comparison among different models. NTN-Diff achieves a comparable inference time while exhibiting the lowest memory cost among all methods.
Methods (Venue) Inference Time (s) $\downarrow$ VRAM Consumption (GB) $\downarrow$
BLD avrahami2023blended (TOG’23) 4.78 2.11
PP zhuang2023task (ECCV’24) 7.24 4.09
BrushNet ju2024brushnet (ECCV’24) 6.59 3.19
HDP manukyan2023hd (ICLR’25) 10.57 11.75
NTN-Diff (Ours) 6.96 2.11
Appendix E Additional Quantitative and Qualitative analysis in the Sec.3.2 of the Mainbody
@@ III. Late Stage
Appendix F More Ablation Studies on the Impact of Text and Null-Text Prompts on Low-and-Mid Frequency Bands during Denoising Process
@@
Appendix G Additional Generation Results for the Ablation Study in Sec.3.3.2 of the mainbody
@@
Appendix H Limitations and Broader Impacts